# Partnering with Generative AI: Experimental Evaluation of Human-Led and Model-Led Interaction in Human-AI Co-Creation

### v1

@@ 6. Discussion
In our research, we aimed to explore human-LLM interaction modes to augment human capabilities in creativity tasks. We focused on the comparison between human-led and model-led interaction formats and examined their influence on the idea quality, idea diversity, and perceived idea ownership. Additionally, we explored the role of perceived cognitive workload, within-subject diversity changes, and participants’ elaboration depth. For this, we conducted a randomized controlled experiment with five different conditions, where participants interacted with a proactive, model-led mode or a reflective, human-led mode (in two variants, i.e., a suggestion-mode and a question-mode). We compared these interaction modes with a vanilla LLM and a control group without LLM support. Our results show that the human-led “flipped” collaboration format—where the LLM asks the participant questions, and the participant is the main creativity contributor—augments human capabilities regarding all three measures. In contrast, providing the participants with suggestions as inspiration does not improve creative outcomes. An exploratory analysis reveals that perceived cognitive workload is higher in the human-led conditions; however, only participants in the question-mode deeply elaborate on the LLM input. Moreover, the question-mode broadens the idea space and supports participants in generating more diverse ideas, whereas the model-led and suggestion-mode conditions lead to homogenization.
@@ 6.1. Question-Mode and Model-Led Improve Idea Quality
Our results show that participants interacting with the question-mode and the model-led collaboration format create ideas with significantly higher idea quality. These findings align with previous research on human-AI co-creation (Lee et al., 2025; Kumar et al., 2025; Holzner et al., 2025; Doshi and Hauser, 2024; Anderson et al., 2024b), which suggests that humans can enhance their creative performance when collaborating with LLMs. However, participants interacting with the suggestion-mode and the vanilla LLM do not create higher-quality ideas, contradicting previous findings (Lee and Chung, 2024; Doshi and Hauser, 2024). One explanation may be task complexity. In simple creativity tasks (such as generating alternative uses for everyday objects like a rope), LLMs can be readily integrated to extend ideas. By contrast, more open-ended and complex tasks place a higher demand on participants and interaction modes. In this context, effective interaction with the vanilla LLM requires a certain prompting experience, as the copy and paste of the task is not sufficient. Also, the usefulness of suggestions may be limited, as participants often need a certain level of domain expertise to evaluate and integrate these suggestions into their ideas.
@@
Another explanation is related to fixation effects (Kohn and Smith, 2011), where users may stick to a certain idea and not develop it further.. For example, the analogies in the suggestion-mode are already provided as a fully articulated example that does not let any room for ambiguity. Such explicitness can constrain human creativity by restricting the solution space (Qin et al., 2025). However, in the question-mode, open-ended questions were presented that allowed ambiguity. This ambiguity allows participants to explore alternative ideas and elaborate on them further (Gaver et al., 2003). In this way, question-mode might be better suited to mitigate the effects of idea fixation compared to suggestions, while, at the same time, still leveraging the generative capabilities of LLMs.
6.2. Questions Broaden Idea Diversity, Suggestions Constrain
@@
Our results show that participants interacting with the question-mode LLM generate significantly more diverse ideas compared to the model-led condition. Our within-subject analysis further highlights that the question-mode not only avoids homogenizing human ideas, but can even reinforce differences: refined ideas after LLM interaction were more diverse than the initial ideas. This suggests that the homogenization effect often attributed to LLMs (Anderson et al., 2024b; Doshi and Hauser, 2024) is not an inherent limitation of the models, but rather an effect of interaction design. Elaborate collaborative frameworks can thus mitigate homogenization and actively foster divergent thinking.
By contrast, the suggestion-mode is overall less effective. Participants who receive suggestions do not produce more diverse ideas than those in the model-led condition. In the within-subject analysis, the ideas are homogenized by the interaction with the suggestion-mode. Our findings extend current literature on the timing of LLM interaction (Qin et al., 2025) by showing that, even when participants first generate their own ideas before seeing LLM suggestions, a homogenizing effect exists. Generating ideas first may reduce, but does not eliminate, fixation, and, crucially, does not lead to augmented creativity.
@@
Interestingly, contrary to Meincke et al. (2025), even the vanilla LLM without any guiding system prompt produced relatively diverse ideas. This indicates that removing strong guidance may broaden the exploration space and counter homogenization. However, the absence of scaffolding also limits opportunities for structured improvement and higher idea quality, suggesting a trade-off between diversity and guidance.
@@
In summary, these results are notable because many existing creativity support tools—both in research and practice—are designed around execution-oriented interactions (similar to our model-led condition) or suggestion-based scaffolds (Doshi and Hauser, 2024; Qin et al., 2025; Shaer et al., 2024). For example, Figma’s ideation Jambot [^2] https://www.figma.com/community/widget/1274481464484630971, or Microsoft’s Copilot in Word [^3] https://support.microsoft.com/en-us/office/welcome-to-copilot-in-word-2135e85f-a467-463b-b2f0-c51a46d625d1, where ideation is scaffolded through suggested content or user instructions are executed in draft form. These tools appear, in particular, to carry a risk of homogenization. Our findings thus support recent claims of the relevance of evaluating creativity support tools via homogenization analysis (Anderson et al., 2024a).
@@ 6.3. Question-Mode Keeps Perceived Idea Ownership High
Our results show that participants interacting with the question-mode perceive significantly higher idea ownership compared to the model-led mode and the vanilla LLM. In line with prior research, high human involvement in the process can increase the perceived idea ownership of participants (Xu et al., 2024). This represents a possibility to effectively co-create with LLMs without diminishing users’ perceived ownership, by keeping them in the lead of the creative work and providing support from the LLM. Thereby, we extend earlier work by showing that not only longer prompts (Joshi and Vogel, 2025) or user-initiated ideation (Qin et al., 2025) but also question-driven interactions can preserve ownership while benefiting from LLM augmentation.
Interestingly, participants in the suggestion-mode do not experience higher ownership, despite being less constrained than in the model-led condition. This contradicts prior findings (Joshi and Vogel, 2025) and may be explained by the lack of deeper elaboration. As our analyses show, participants in the suggestion-mode often fail to integrate suggestions with their own ideas, which undermines their sense of contribution.
In the question-mode, participants are thus effectively supported by an LLM, but with still high perceived idea ownership. Our findings show that, to augment human capabilities with LLMs and mitigate negative effects like reduced idea diversity and idea ownership, it is effective to place the human in the creative lead. When designing LLMs as a thought partner, ideas can emerge through the interaction, augmenting human creativity rather than replacing it.
@@ 6.5. Practical Implications for the Design of Co-Creation Frameworks
Although LLMs are often described as thought partners with an emphasis on co-creation as a design pattern (Rafner et al., 2023), actual design for LLM-based co-creation remains scarce. Our study thus provides designers of co-creation tools with three concrete design principles on how to effectively create co-creation interactions with LLMs. Based on our findings, we suggest that designers of co-creation tools should:
  • Principle 1: Create with users, not for them: LLM interaction modes should be designed to augment human cognitive processes rather than replace them. In creative work, reducing human effort may accelerate creative processes but risks quality, and perceived idea ownership. Effective co-creation resembles good mentorship: supporting the process of exploration and discovery rather than providing ready-made answers.
  • Principle 2: Make the user better, not just the system: Our results show that the most effective collaboration format might not be the one with the most capable technology, but the one that best supported human thinking. Designing for creativity thus requires a human-centered approach that identifies where users benefit most from support and structures AI contributions to foster synergy.
  • Principle 3: Provide users with guidance: LLMs are compelling because of their general-purpose approach. However, this openness might lead to uncertainty and frustration for the user, as they might not be sure which tasks they can delegate to the AI and where they should choose a certain collaboration format. This presents an opportunity for co-creation tool design: by guiding users toward appropriate interaction formats for specific tasks, systems can help people leverage AI effectively.
6.6. Limitations and Outlook
@@
Despite these constraints, our results highlight the augmenting potential of question-mode interaction in creativity contexts. Future studies should test whether this interaction mode can also mitigate risks of reduced critical thinking (Lee et al., 2025) or diminished skill acquisition (Kumar et al., 2025). A “flipped” collaboration format—in which humans remain in the lead while the LLM supports reflection—may help ensure that skills persist beyond individual interactions. This raises the broader question of whether LLMs should always do everything they are capable of. In the context of co-creation, the most effective LLM is the one that best supports the user in their creativity. Yet, many applications continue to struggle with implementing LLMs in ways that deliver meaningful benefits. While more advanced models may improve outcomes, in co-creation, they constitute only one side of the equation. The other side—the design of interaction modes—still bears substantial, and largely untapped, potential.
@@ Acknowledgements.
Appendix A Task Description and Full Questionnaire
Task Description
Come up with a creative way to repurpose existing car features—such as cameras, sensors, screens, lights, or speakers—for entirely new uses beyond their original purpose in vehicles. This idea can be applied inside the vehicle in new ways or outside of the automotive context altogether. Your goal: Reimagine how this technology could serve new functions, solve problems, or create novel experiences in completely different settings. Important: Please do not use AI like ChatGPT at this point of the task. We want to see your own creative ideas.
Questionnaire Items and Response Scales
Creative Self-Efficacy
Scale: 1 = Strongly Disagree, 7 = Strongly Agree
  1. (1)
    I was good at coming up with novel ways to repurpose vehicle features.
  2. (2)
    I had confidence in my ability to solve problems creatively while working on the creativity task.
  3. (3)
    To ensure you are reading carefully, please select Option 2 for this question.
  4. (4)
    I had a knack for further developing the ideas in this exercise.
Perceived Technology Agency
Scale: 1 = Strongly Disagree, 5 = Strongly Agree
  1. (1)
    The chatbot goes its own way.
  2. (2)
    The chatbot does things by itself.
  3. (3)
    The chatbot takes initiative.
  4. (4)
    The user of this chatbot does not have to bother about this chatbot a lot.
  5. (5)
    The chatbot works independently.
NASA-TLX (Workload)
Scale: 0 = very low, 100 = very high (slider)
  1. (1)
    How mentally demanding was the task?
  2. (2)
    How hard did you have to work to accomplish your level of performance?
Performance Expectancy
Scale: 1 = Strongly Disagree, 7 = Strongly Agree
  1. (1)
    I find the chatbot useful for creative tasks.
  2. (2)
    Using the chatbot helps me arrive at creative ideas more quickly.
  3. (3)
    Using the chatbot increases my productivity when I am working on creative tasks.
Effort Expectancy
Scale: 1 = Strongly Disagree, 7 = Strongly Agree
  1. (1)
    Overall, I believe the chatbot is easy to use.
  2. (2)
    Learning how to use the chatbot for the creativity task is easy for me.
  3. (3)
    My interactions with the chatbot are clear and understandable.
  4. (4)
    It is easy for me to become skillful at using the chatbot for creativity tasks.
Hedonic Motivation
Scale: 1 = Strongly Disagree, 7 = Strongly Agree
  1. (1)
    I think using the chatbot is fun.
  2. (2)
    I think using the chatbot is very entertaining.
  3. (3)
    I think using the chatbot is enjoyable.
AI Experience
Scale: 1 = Never, 2 = Rarely, 3 = Sometimes, 4 = Often, 5 = Daily
  1. (1)
    How often do you use AI systems like ChatGPT during the week?
Propensity to Trust Technology
Scale: 1 = Strongly Disagree, 5 = Strongly Agree
  1. (1)
    I usually trust AI-powered chatbots until there is a reason not to.
  2. (2)
    For the most part, I distrust AI-powered chatbots.
  3. (3)
    In general, I would rely on an AI-powered chatbot to assist me.
  4. (4)
    This is an attention check question. Please select Strongly Agree (option 5) to show you are paying attention.
  5. (5)
    My tendency to trust AI-powered chatbots is high.
  6. (6)
    It is easy for me to trust AI-powered chatbots to do their job.
  7. (7)
    I am likely to trust AI-powered chatbots even when I have little knowledge about it.
Perceived Idea Ownership
Scale: 1 = Strongly Disagree, 7 = Strongly Agree
  1. (1)
    I feel like I am the creator of the idea.
  2. (2)
    I am the main contributor to the content of the resulting idea.
  3. (3)
    I have made substantial contributions to the content of the resulting idea.
  4. (4)
    I am accountable for all aspects of the idea.
  5. (5)
    I am responsible for at least part of the idea.
Appendix B Prompts Used in the Different Conditions
Table 4. Experimental Conditions and AI Assistant Prompts
@@ Appendix B Prompts Used in the Different Conditions
Appendix C Examples of Ideas

### v2

@@ 7. Discussion
In our research, we aimed to explore human-LLM interaction modes to augment human capabilities in creativity tasks. We focused on the comparison between human-led and model-led interaction formats and examined their influence on the idea quality, idea diversity, and perceived idea ownership. Additionally, we explored the role of perceived cognitive workload, within-subject diversity changes, and participants’ elaboration depth. For this, we conducted a randomized controlled experiment with five different conditions, where participants interacted with a proactive, model-led mode or a reflective, human-led mode (in two variants, i.e., a suggestion-mode and a question-mode). We compared these interaction modes with a vanilla LLM and a control group without LLM support. Our results show that the human-led “flipped” collaboration format—where the LLM asks the participant questions, and the participant is the main creativity contributor—augments human capabilities regarding all three measures. In contrast, providing the participants with suggestions as inspiration does not improve creative outcomes as expected. An exploratory analysis reveals that perceived cognitive workload is higher in the human-led conditions; however, only participants in the question-mode deeply elaborate on the LLM input. Moreover, the question-mode broadens the idea space and supports participants in generating more diverse ideas, whereas the model-led and suggestion-mode conditions lead to homogenization. Overall, our findings are robust across different domains, tasks, and creativity measures (see the validation study in Section 6).
@@ 7.1. Question-Mode and Model-Led Improve Idea Quality
Across both experiments, our results show that participants interacting with the question-mode and the model-led collaboration format create ideas with significantly higher idea quality. These findings align with previous research on human-AI co-creation (Lee et al., 2025; Kumar et al., 2025; Holzner et al., 2025; Doshi and Hauser, 2024; Anderson et al., 2024b), which suggests that humans can enhance their creative performance when collaborating with LLMs. However, participants interacting with the suggestion-mode and the vanilla LLM do not create higher-quality ideas, contradicting previous findings (Lee and Chung, 2024; Doshi and Hauser, 2024). One explanation may be task complexity. In simple creativity tasks (such as generating alternative uses for everyday objects like a rope), LLMs can be readily integrated to extend ideas. By contrast, more open-ended and complex tasks place a higher demand on participants and interaction modes. In this context, effective interaction with the vanilla LLM requires a certain prompting experience, as the copy and paste of the task is not sufficient. Also, the usefulness of suggestions may be limited, as participants often need a certain level of domain expertise to evaluate and integrate these suggestions into their ideas.
@@
Another explanation is related to fixation effects (Kohn and Smith, 2011), where users may stick to a certain idea and not develop it further. For example, the analogies in the suggestion-mode are already provided as a fully articulated example that does not let any room for ambiguity. Such explicitness can constrain human creativity by restricting the solution space (Qin et al., 2025). However, in the question-mode, open-ended questions were presented in a way that allowed ambiguity. This ambiguity allows participants to explore alternative ideas and elaborate on them further (Gaver et al., 2003). In this way, the question-mode might be better suited to mitigate the effects of idea fixation compared to suggestions, while, at the same time, still leveraging the generative capabilities of LLMs.
Notably, the exploratory analysis in the context of product innovation for UK students revealed that the higher idea quality observed in the question-mode and model-led (one-shot) conditions translated to descriptively higher purchase intent ratings from the target group. This extends previous studies conducted by Girotra et al. (2023), thereby demonstrating that LLMs can generate commercially viable product ideas while maintaining idea diversity through appropriate interaction design.
7.2. Questions Broaden Idea Diversity, Suggestions Can Constrain
@@
Our results consistently show that participants interacting with the question-mode LLM generate significantly more diverse ideas compared to the model-led one-shot condition. Our within-subject analysis further highlights that the question-mode not only avoids homogenizing human ideas, but can even reinforce differences: refined ideas after LLM interaction were more diverse than the initial ideas. This suggests that the homogenization effect often attributed to LLMs (Anderson et al., 2024b; Doshi and Hauser, 2024) is not an inherent limitation of the models, but rather an effect of interaction design. Elaborate collaborative frameworks can thus mitigate homogenization and actively foster divergent thinking.
By contrast, the effectiveness of the suggestion-mode in supporting idea diversity might depend on the domain. Whereas in the initial AUT experiment, the suggestion-mode rather constrained the idea space, it kept idea diversity high in the validation study. A reason for that might be, that the AUT task in the first study did not inspire deeper elaboration in the suggestion-mode. Similar to Kumar et al. (2025), participants might have simply adopted LLM suggestions. In contrast, creating product ideas requires participants to identify genuine needs and problems, enabling the suggestions as inspiration rather restriction, thereby fostering idea diversity. Our findings extend current literature on the homogenization effect by showing that not only the timing of LLM support (Qin et al., 2025) or prompting matters (Meincke et al., 2024b), but also the specific interaction mode and the task context.
@@
Notably, contrary to Meincke et al. (2025), even the vanilla LLM without any guiding system prompt produced relatively diverse ideas. This might be partly explained by participants generating ideas that did not strictly adhere to task constraints. Additionally, this indicates that removing strong guidance may broaden the exploration space and counter homogenization. However, the absence of scaffolding also limits opportunities for structured improvement and higher idea quality, suggesting a trade-off between diversity and guidance. Interestingly, participants in the iterative model-led condition of the validation study did not exhibit homogenization, underscoring the value of iterative processes in creative work (Wallas, 1926; Guilford, 1967; Frich et al., 2019).
@@
In summary, these results are notable because many existing creativity support tools—both in research and practice—are designed around execution-oriented interactions (similar to our one-shot model-led condition) or suggestion-based scaffolds (Doshi and Hauser, 2024; Qin et al., 2025; Shaer et al., 2024). For example, Figma’s ideation Jambot [^3] https://www.figma.com/community/widget/1274481464484630971, or Microsoft’s Copilot in Word [^4] https://support.microsoft.com/en-us/office/welcome-to-copilot-in-word-2135e85f-a467-463b-b2f0-c51a46d625d1, where ideation is scaffolded through suggested content or user instructions are executed in draft form. These tools appear, in particular, to carry a risk of homogenization. Our findings thus support recent claims of the relevance of evaluating creativity support tools via homogenization analysis (Anderson et al., 2024a).
@@ 7.3. Question-Mode Keeps Perceived Idea Ownership High
Across both experiments, our results show that participants interacting with the question-mode perceive significantly higher idea ownership compared to the model-led mode and the vanilla LLM. In line with prior research, high human involvement in the process can increase the perceived idea ownership of participants (Xu et al., 2024). This represents a possibility to effectively co-create with LLMs without diminishing users’ perceived ownership, by keeping them in the lead of the creative work and providing support from the LLM. Thereby, we extend earlier work by showing that not only longer prompts (Joshi and Vogel, 2025) or user-initiated ideation (Qin et al., 2025) but also question-driven interactions can preserve ownership while benefiting from LLM augmentation. Interestingly, participants in the suggestion-mode only experience higher ownership in the second study in the domain of product ideation, but not in the first study using an AUT task, mirroring the pattern observed for idea diversity.
In summary, in the human-led conditions—particularly in the question-mode—participants were effectively supported by the LLM while maintaining high perceived idea ownership. Our findings show that, to augment human capabilities with LLMs and mitigate negative effects like reduced idea diversity and idea ownership, it is effective to place the human in the creative lead. When designing LLMs as a thought partner, ideas can emerge through the interaction, augmenting human creativity rather than replacing it.
@@ 7.5. Practical Implications for the Design of Co-Creation Frameworks
Although LLMs are often described as thought partners with an emphasis on co-creation as a design pattern (Rafner et al., 2023; Collins et al., 2024), actual design for LLM-based co-creation remains scarce. Our study provides designers of co-creation tools with three concrete design principles on how to effectively create co-creation interactions with LLMs. Based on our findings, we suggest that designers of co-creation tools should:
  • Principle 1: Maintain human exploration of the idea space: To integrate LLMs in co-creation applications, HCI designers should amplify human cognitive exploration, allowing users to navigate broad possibility spaces with creativity and autonomy. Our findings imply that, in creative work, a model-led interaction approach may accelerate creative processes but at the risk of diversity and perceived idea ownership. This is particularly salient during exploration stages, where human creativity offers unique strengths in navigating broad possibility spaces. In practice, this means that the human initially generates ideas, while the LLM primarily asks questions that prompt elaboration and reflection. This helps establish a mutual understanding between human and AI regarding the intentions, a key principle for effective co-creation (Collins et al., 2024).
  • Principle 2: Counter cognitive offloading: The most effective LLM-based co-creation tools will not necessarily rely on the most capable technology, but on designs that actively support human thinking. Co-creation with model-led LLM interaction modes can lead to overreliance and cognitive offloading. This is not only important for the exploration stage, but also for evaluating the best outcome, as only an individual engaging with the output can evaluate it properly (Tankelevitch et al., 2024; Gmeiner et al., 2025). Our results suggest that designing for creativity thus requires a human-centered approach that engages humans in deep elaboration. In practice, designers should therefore aim to promote active human contribution rather than passive generation of LLM output.
  • Principle 3: Guide users through iterative refinement with LLMs: Building on well-established theories of iterative creativity (Resnick et al., 2005), LLM systems for co-creation should not treat generation as a one-shot process but as an evolving collaboration between user and model. While LLMs can rapidly generate output in a single step, this approach often fails to produce optimal and diverse results. Moving forward, LLM-based co-creation tools should intentionally scaffold iterative refinement—whether led by the user or by the model—to help users explore alternatives, critique drafts, and converge on stronger solutions. Designing for this iterative loop implies that interfaces should support revision pathways (e.g., models could learn to propose structured next steps), and systems should encourage reflection and re-generation. Embedding iteration as a core interaction pattern will unlock greater creativity, higher-quality ideas, and overall more effective human–AI co-creation.
These principles can be, for example, applied to LLM-based writing assistants for creative writing (e.g., for essays or short stories) as Microsoft’s Copilot in Word. The LLM assistant can first support idea generation through a brief question-only stage that prompts users to articulate claims, audiences, and constraints, keeping early exploration human-led. During elaboration, the assistant can encourage users to outline key points before any drafting occurs, reducing cognitive offloading and preserving meaningful human contribution. During refinement, initiative might be shared between human and LLM, with the human shaping the text and the AI providing feedback and additional ideas across iterative cycles.
7.6. Limitations and Potential for Future Work
@@
Our study examined text-based divergent thinking tasks with general population participants in controlled conditions. While this design enabled rigorous hypothesis testing, this points to concrete opportunities for extending our work. Nevertheless, we validated our findings in a more realistic ideation task (see Section 6). Future research should build on our findings by conducting field studies in authentic work settings, testing interaction modes with domain experts, and examining how different task contexts moderate the effectiveness of human-led versus model-led approaches. First, field studies would strengthen external validity. Deploying these interaction modes in real-world creative work—such as marketing professionals developing campaigns—would test whether our laboratory findings generalize to authentic practice where creativity must balance strategic constraints (brand consistency, search engine advertising requirements, character limits). Such studies would also enable measuring real-world success metrics like conversion rates beyond creative quality ratings. Second, testing with domain experts is essential. Research suggests that experts might not only be better at generating creative ideas but also at evaluating the novelty and relevance of AI-generated outputs (Eisenmann et al., 2025; Jia et al., 2024), potentially affecting the effectiveness of different interaction modes. Understanding whether expertise moderates interaction mode effectiveness would clarify the boundary conditions of our findings. Finally, different task domains have different requirements. While question-mode interactions support divergent thinking, they might hinder performance in programming or problem-solving where productivity and correctness outweigh novelty. Future studies should examine interaction modes across domains and explore approaches that adaptively support human strengths (Collins et al., 2024).
Beyond these specific directions, our results highlight the augmenting potential of question-mode interaction in creativity contexts. Future studies should test whether this interaction mode can also mitigate risks of reduced critical thinking (Lee et al., 2025) or diminished skill acquisition (Kumar et al., 2025). A “flipped” collaboration format—in which humans remain in the lead while the LLM supports reflection—may help ensure that skills persist beyond individual interactions. This raises the broader question of whether LLMs should always do everything they are capable of. In the context of co-creation, the most effective LLM is the one that best supports the user in their creativity. Yet, many applications continue to struggle with implementing LLMs in ways that deliver meaningful benefits. While more advanced models may improve outcomes, in co-creation, they constitute only one side of the equation. The other side—the design of interaction modes—still bears substantial, and largely untapped, potential.
@@ Acknowledgements.
Data availability
The questionnaire, analysis code, and anonymized data are publicly available in a GitHub repository: https://github.com/SM2982/replication-creativity-CHI2026.
Appendix A Prompts Used in the Different Conditions
Table 5. Experimental Conditions and AI Assistant Prompts for the Initial Study
@@ Appendix A Prompts Used in the Different Conditions
Table 6. Experimental Conditions and AI Assistant Prompts for the Validation Study
Condition Prompt description and key instructions
Feedback (Question-mode) Exploration Phase:¡Your task:¿ You act as a creative thought partner. The user has shared an idea and your goal is to help the user explore unexpected and novel ideas for the task [product ideas for UK university students]. Ask the user a thought-provoking question, to generate alternative ideas to the initial idea. ¡The questions you ask¿ 1. For the first alternative idea: Using the knowledge approach: To help them reimagine or enhance this idea, briefly share 1–2 sentences of specific knowledge about: Specific product features, other products or market trends that could contribute to or transform the product of the initial idea. Make clear that this knowledge could inspire a new or improved version of the idea. Then ask: ’How could this insight inspire a new version or twist on your idea? Please propose an alternative idea.’ 2. For the second alternative idea: Using the analogy approach: Shift to a more lateral thinking approach. Provide a short analogy—this could be a product, brand, use case, or concept unrelated to the initial idea, but that could inspire a new version or twist on the initial idea. Use 1–2 sentences to describe the analogy clearly. Then ask: ’How could this analogy inspire a new version or twist on your idea? Please propose an alternative idea.’ ¡How you act¿ Shortly acknowledge the user input. Ask thought-provoking questions using the knowledge approach or the analogy approach. Sound like smooth conversation with a friend. Max 40 words. Only ask one question per message. Write relevant words in **bold** using markdown formatting.”
Choice Phase:¡Your task:¿ Present the alternatives clearly to the user and help them choose which one to develop further. Present all available options verbatim without further elaboration (initial idea + alternatives). Write relevant words in **bold** using markdown formatting. Ask the user ’Please choose the most **unique and creative** option to develop further’. When the user makes their choice, call the choose_alternative function.”
Elaboration Phase:¡Your task:¿ You are an expert in facilitating practical idea development. The user has shared a creative product idea for university students in the UK. Your goal is to help them further elaborate and clarify the idea—especially when it is still abstract or vague—so it becomes more useful and actionable. Briefly acknowledge the user’s previous message. Ask one focused, thought-provoking question that helps the user: Ask a question that might further increase the creativity and novelty of the idea, Clarify what problem the idea solves for UK university students, Specify the context of use or the intended users, Make abstract aspects of the idea more concrete, Consider the specific needs, behaviors, habits, or preferences of UK university students and their usage scenarios, Shift perspective (e.g., different user groups, environments) to inspire elaboration, Identify what may be missing in the current idea (e.g., unclear value, user role, context) and ask a targeted question to address this gap. If the idea is vague or confusing, say so and ask a specific question to clarify it. Be critical: if something is hard to understand or doesn’t seem feasible, say that clearly. Use markdown formatting **bold** to highlight key concepts or terms from the user’s idea.”
Suggestion-mode Exploration Phase: “You are an expert in facilitating creative idea development through suggestions. The user has shared an idea for the task [product ideas for UK university students] and your goal is to help them explore it more deeply and make it more novel. Provide 3 specific suggestions for creative ideas (40-60 words each) that are similar to the user’s idea but think it further and make it more novel and surprising. Your suggestions should focus on: Combining the idea with unrelated contexts, Increasing uniqueness and originality, Creating unexpected elements, Challenging assumptions, Reversing or inverting core concepts, use first principles thinking and lateral thinking, Prioritize novelty, unexpected connections and contrasting perspectives over safe ideas. Format: *Some ideas to make your concept more creative and novel:* **Suggestion 1:** [novel suggestion] **Suggestion 2:** [novel suggestion] **Suggestion 3:** [novel suggestion].”
Elaboration Phase: “You are an expert in facilitating practical idea development through suggestions. The user has shared a creative idea for the task [product ideas for UK university students] and your goal is to help them make it more useful and elaborate. Provide 3 specific suggestions (40-60 words each) that enhance the user’s idea to make it more useful. Your suggestions should focus on: Solving specific problems or addressing frustrations of UK university students, Identifying clear target users and benefits of UK university students, Enhancing the value proposition, Addressing implementation challenges of UK university students, Creating competitive advantages, Improving commercial potential and monetization, The three suggestions should be very different from each other. Format: Some ideas to make your concept more useful and practical: **Suggestion 1:** [usefulness-focused suggestion] **Suggestion 2:** [usefulness-focused suggestion] **Suggestion 3:** [usefulness-focused suggestion].”
Model-led (one-shot) System Prompt:¡Your task¿ Your task is to take the user’s input and write an idea by the following criteria: (1) It has to fit the task [product ideas for UK university students] and the product idea context for university students in the UK (2) Increase the novelty and creativity of the idea (3) Increase the usefulness of the idea (4) Make it more elaborate (5) Provide one coherent idea. ¡Your Response¿ Keep your response focused on the improved idea - don’t add explanations, disclaimers, or ask follow-up questions. Please keep the style and the tone of the original idea and keep it as concise as possible, maximum 40 words.”
Model-led (iterative) Exploration Phase: “You are an expert in facilitating creative idea development through suggestions. The user has shared an idea for the task [product ideas for UK university students] and your goal is to help them explore it more deeply and make it more novel. Provide 3 specific alternative creative ideas (40-60 words each) that are inspired by the user’s idea but think it further and make it more novel and surprising. Your suggestions should focus on: Combining the idea with unrelated contexts, Increasing uniqueness and originality, Creating unexpected elements, Challenging assumptions, Reversing or inverting core concepts, use first principles thinking and lateral thinking, Prioritize novelty, unexpected connections and contrasting perspectives over safe ideas, Each suggestion must be self-contained and readable as a standalone proposition. Format: *Some ideas to make your concept more creative and novel:* **Suggestion 1:** [novel suggestion] **Suggestion 2:** [novel suggestion] **Suggestion 3:** [novel suggestion]. After the user selects and refines one of these suggestions, automatically move to the next stage.”
Elaboration Phase: “You are an expert in facilitating practical idea development through suggestions. The user has shared a creative idea for the task [product ideas for UK university students] and your goal is to help them make it more useful and elaborate. Provide 2 specific suggestions for creative ideas (40-60 words each) that are similar to the user’s idea but think it further and make it more novel and surprising. Your suggestions should focus on: Solving specific problems or addressing frustrations of UK university students, Identifying clear target users and benefits of UK university students, Enhancing the value proposition, Addressing implementation challenges of UK university students, Creating competitive advantages, Improving commercial potential and monetization, The two suggestions should be very different from each other, Each suggestion must be self-contained and readable as a standalone proposition. Format: Some ideas to make your concept more useful and practical: **Suggestion 1 (Your Refined Idea):** [user’s refined idea] **Suggestion 2:** [usefulness-focused suggestion] **Suggestion 3:** [usefulness-focused suggestion].”
Appendix B Examples of Ideas (Initial Study)